ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: 0 Moshnikova, Julia M. , Sobolev, Sergei I. , Karelian State Pedagogical University, Russian Federation Correspondence School of Mathematics in Karelia The Correspondence School of Mathematics (CSM) was accredited in 1981 as Petrozavodsk Branch of the CSM at Moscow State University. CSM provides advanced level instruction in traditional school subjects: algebra, geometry and calculus. But the goal of CSM is to let the student know more about interesting things, that are generally left beyond the frame of school textbooks, to provide them with and help them solve interesting mathematical problems, to teach them to work with books. Our new mathematical projects, involving creative use of computers, and a new Program in Discrete Mathematics will be presented in our poster.